40% of all workers of an industry are college graduates. If a sample of 200 workers is randomly selected, what is the probability that a college graduate in the sample would be

A. more than 80?

B. between 45% and 48%?



1
Expert's answer
2022-01-04T17:50:47-0500

Solution:

"p=40\%=0.4,q=0.6,n=200
\\ X\sim Bin(n,p)"

"X\sim N(\mu,\sigma)"

"\mu=np=200(0.4)=80
\\ \sigma=\sqrt{npq}=\sqrt{200(0.4)(0.6)}=\sqrt{48}=6.93"

A. the probability that a college graduate in the sample would be more than 80

"=P(X>80)
\\=P(X>80.5)" [using continuity correction]

"=1-P(X\le 80.5)
\\=1-P(Z\le \dfrac{80.5-80}{6.93})
\\=1-P(Z\le 0.072)
\\=1-0.52790
\\=0.4721"


B. 45% of 200 = 90, 48% of 200 = 96

the probability that a college graduate in the sample would be between 45% and 48%

"=P(90<X<96)
\\=P(X<96)-P(X<90)
\\=P(X<95.5)-P(X<89.5)" [using continuity correction]

"=P(Z<\dfrac{95.5-80}{6.93})-P(Z<\dfrac{89.5-80}{6.93})
\\=P(Z<2.24)-P(Z<1.37)
\\=0.98745-0.91466
\\=0.07279"